# A Frank-Wolfe-based primal heuristic for quadratic mixed-integer optimization

### v1

@@ 5.5 Impact of the Frank-Wolfe accuracy
We aim to strike a balance between computational effort in the convex relaxations and solution accuracy. Too few iterations would cause the relaxation solutions not to move sufficiently from one node to the next, while too many iterations would make the framework spend too much time computing high-accuracy solutions, which will in any case only serve as starting point to compute integer-feasible solutions through e.g., rounding afterwards. Our experiments show that 10 to 100 FW iterations per node yields a good balance between accuracy and relaxation cost, in terms of number of instances with at least one solution, mean residual gap, and primal integral.
@@
Table 6: Performance of the base heuristic with different power penalty parameters $p$. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
@@
5.7 Competition and Improvements on QPLIB
Table 7: Improved solutions for QPLIB instances.
Instance Obj. Sense   Previous Best   New Solution    Gap
QPLIB_2169 max 29.0 30.0 3.45
QPLIB_2174 max 150.0 152.0 1.33
QPLIB_2205 max 88.0 90.0 2.27
QPLIB_3347 min 3,819,920 3,818,879 0.03
QPLIB_3584 min -25,254 -25,386 0.52
QPLIB_3709 min 5,726,530 5,710,645 0.28
QPLIB_3860 min -19,685 -20,161 2.42
QPLIB_10022 min 6,267,782 1,374,066 78.07

### v2

@@ 6.5 Impact of the Frank-Wolfe accuracy
We aim to strike a balance between computational effort in the convex relaxations and solution accuracy. Too few iterations would cause the relaxation solutions not to move sufficiently from one node to the next, while too many iterations would make the framework spend too much time computing high-accuracy solutions, which will in any case only serve as starting point to compute integer-feasible solutions through, e.g., rounding afterwards. Our experiments show that 10 to 100 FW iterations per node yields a good balance between accuracy and relaxation cost, in terms of number of instances with at least one solution, mean residual gap, and primal integral.
@@
Table 8: Performance of the base heuristic with different power penalty parameters pp. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
@@ 7 Conclusion
Acknowledgements
We thank the organizing committee of the Land-Doig MIP Computational Competition, with special thanks to Aleksandr Kazachkov and Jan Kronqvist. Research reported in this paper was partially supported through the Research Campus Modal funded by the German Federal Ministry of Research, Technology, and Space (fund numbers 05M14ZAM, 05M20ZBM) and the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+. The work of Mathieu Besançon benefited from the support of the FMJH Program PGMO and from MIAI at Université Grenoble Alpes (grant ANR-19-P3IA-0003).